OpenStudy (anonymous):

Consider the problem: “The perimeter of a rectangle is more than 24 cm. If the length of the rectangle is 3 more than the width, what are the minimum dimensions of the rectangle?” What are the quantities to be determined in the problem?

OpenStudy (campbell_st):

let x be the width... then perimeter is \[P = 2 \times (3x) + 2\times (x)\] P = 8x using the inequality 8x > 24 then x > 3 so the width is greater than 3

OpenStudy (anonymous):

What are the quantities to be determined in the problem?

OpenStudy (campbell_st):

the quantities are the width must be greater than 3 cm.... which also means the length will be greater than 9cm

OpenStudy (anonymous):

Okay, thank you :)

OpenStudy (mertsj):